what are 2 numbers that would make an average of 92

Found 2 solutions by Alan3354, macston:
Answer by Alan3354(69443)

Any 2 numbers with a sum of 184.
n1 = 184 - n2
Pick any number for n2, find n1.

Answer by macston(5194)

.
F=first number; S=second number
.
F+S=2(92)
F+S=184
S=184-F
.
Pick any number (first number), F.
Subtract F from 184 to get S.
F and S are two numbers that average 92.
EXAMPLE:
Pick the number 10.
First number is 10.
Second number=184-first=184-10=174
The two numbers are 10 and 174
.
Find the average:
(10+174)/2=92
184/2=92
92=92